In 2008, 58% of adults in a certain country had a college education. A researcher suspects that the percent is higher now. She obtains a random sample of 1012 adults and finds that 590 have a college education. Is this sufficient evidence to support the researcher’s claim at the 10% level of significance? Your answer should show all four steps of the hypothesis test.
Answer #1

Solution :

This is the right tailed test .

The null and alternative hypothesis is

H0 : p = 0.58

Ha : p > 0.58

\hat p = x / n = 590 / 1012 = 0.583

P0 = 0.58

1 - P0 = 1 - 0.58 = 0.42

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.583 - 0.58 / [\sqrt(0.58 * 0.42) / 1012]

= 0.19

P(z > 0.19) = 1 - P(z < 0.19) = 1 - 0.5753

P-value = 0.4247

\alpha = 0.10

P-value > \alpha

Fail to reject the null hypothesis .

There is not sufficient evidence to support the researcher’s claim